# DAPPER: Discriminability-Aware Policy-to-Policy Preference-Based Reinforcement Learning for Query-Efficient Robot Skill Acquisition

### v1

@@ V-E2 Results
V-F Scalability with Respect to Feature Dimensions
V-F1 Settings
V-F2 Results
V-G Real-World Application
In this section, we apply the learned control policies from simulation to real-world environments using a sim-to-real transfer approach [16]. The transfer policy is trained in a domain-randomized environment using the preference reward model $R^{H}$ obtained from human queries. The randomized parameters follow the same settings as in previous work [16]. As shown in Fig. 7, the learned policies were applied in the real world. For details of the robot motion, please refer to the supplementary videos.
@@ VI Discussion
This paper focuses on preference discriminability and proposes a robust learning framework that performs effectively even under high discriminability thresholds (i.e., “Large”). However, in practice, some queries that fall within these thresholds may not be labeled as indistinguishable by human annotators but instead receive incorrect superiority labels as shown in Fig. 3(c). This type of labeling noise has been addressed in several prior works [3, 4]. Integrating such noise-robust methods with our proposed approach presents a promising direction for future research.
@@
Some consideration is necessary when defining the terms of feature extraction [20]. In particular, if multiple similar features are selected and their optimization objectives happen to conflict, it may hinder policy learning. For example, if satisfying different features requires the robot to exhibit opposing behaviors, achieving the target may become difficult or even infeasible. While such cases are unlikely in well-structured tasks, they remain a possibility, and attention to feature design can help avoid such issues.
@@
In our framework, policy constraints (Table II) are used to enforce basic requirements such as movement speed or safety, rather than to represent human preferences. These constraints do not need to be strictly maximized, but only sufficiently satisfied within a margin. This approach is particularly effective because such low-level objectives are easy to define and help improve query discriminability. In contrast, high-level preferences such as motion style are difficult to specify explicitly, and are therefore learned through PbRL. This separation allows the system to delegate well-defined tasks to constraints, while using PbRL to capture complex, preference-based behaviors.

### v2

@@ VI Discussion
Robustness to Label Noise in Preference Annotations: This paper focuses on preference discriminability and proposes a robust learning framework that performs effectively even under high discriminability thresholds (i.e., “Large”). In practice, however, some queries within these thresholds may not be labeled as indistinguishable by human annotators but instead receive incorrect superiority labels, as shown in Fig. 3(c). In addition, as shown in Table IV, we observed that the proposed framework can still operate effectively even when the generated labels include a certain amount of mislabeling (approximately $15\text{\,}\%$), suggesting robustness to imperfect annotations. This type of labeling noise has also been studied extensively in several prior works [8, 9]. Building on these insights, integrating such noise-robust methods with our framework presents a promising direction for future research.
@@
Sensitivity to Hand-crafted Feature Design and Future Direction Toward Automated Extraction: Some consideration is necessary when defining the terms of feature extraction [18]. In particular, the success of our framework still relies on a set of carefully designed features, and the overall performance is sensitive to the quality of this feature selection. If multiple similar features are selected and their optimization objectives happen to conflict, it may hinder policy learning. For example, if satisfying different features requires the robot to exhibit opposing behaviors, achieving the target may become difficult or even infeasible. Likewise, if the selected features are noisy or fail to capture informative aspects of user preferences, the learned policy can become suboptimal. While such cases are unlikely in well-structured tasks, they remain a possibility, and attention to feature design is crucial. A promising future direction is to reduce this dependence on hand-crafted features by integrating automated feature extraction or richer representations learned from high-dimensional sensor data.
Trade-off Between Computational Cost and Policy Diversity: An additional aspect to consider is the computational cost associated with retraining a new policy from scratch after each reward update in order to learn policies with different feature representations. While this approach is more expensive than continuing from previously trained policies, it was adopted to avoid policy bias and to ensure diversity among the learned policies. This highlights a trade-off between computational efficiency and policy diversity: retraining increases training time but provides more reliable exploration and reduces the risk of converging to biased behaviors. We recognize that this trade-off becomes a critical concern when extending the framework to more complex and computationally demanding domains, such as large-scale decision-making problems that require semantic or linguistic understanding. To mitigate this limitation, potential directions include warm-starting policy initialization within a range that avoids policy bias, or leveraging meta-learning techniques to initialize policies in a cost-effective manner. These strategies could help reduce computational overhead while maintaining sufficient diversity across learned policies, making the framework more scalable to large-scale or high-dimensional applications.
@@
Balancing Low-level Constraints and High-level Preferences: In our framework, policy constraints (Table II) are used to enforce basic requirements such as movement speed or safety, rather than to represent human preferences. These constraints do not need to be strictly maximized, but only sufficiently satisfied within a margin. This approach is particularly effective because such low-level objectives are easy to define and help improve query discriminability. In contrast, high-level preferences such as motion style are difficult to specify explicitly, and are therefore learned through PbRL. This separation allows the system to delegate well-defined tasks to constraints, while using PbRL to capture complex, preference-based behaviors.
Stability of the Discriminability Reward During Early Training: A potential concern in our framework is the stability of the discriminability reward $R^{D}$ during the early stages of training. Since the discriminator is initially trained on only a small number of labels, its preference discriminability estimates may be noisy or biased, which could misguide policy exploration and lead to suboptimal behaviors. This sensitivity can become more noticeable as the input and output dimensionality of the discriminator increases, making it harder to obtain reliable preference discriminability scores from limited data. In our experiments, however, such instability was not severe, partly because the feature dimensionality was relatively small and the discriminator $D$ incorporated Monte Carlo dropout [16], which helped mitigate instability in its outputs. Nevertheless, instability remains a possible challenge in more complex scenarios, and developing mechanisms to further stabilize the $R^{D}$ represents an important direction for future work, for example by incorporating dimensionality reduction techniques or attention mechanisms.
Toward Capturing Richer and Higher-dimensional Human Preferences: While our framework takes a step toward making PbRL more practical by explicitly addressing query discriminability, it still falls short of fully capturing human preferences. The feature dimensionality in our experiments (up to six) may overlook subtle aspects of human preference, such as smoothness, energy efficiency, or style. Moreover, learning a satisfactory policy still requires hundreds of queries, although Fig. 4 shows that our framework reduces this number to about one-tenth of that required by previous approaches. These findings indicate that further progress toward real human experiments will require richer, higher-dimensional preference features and additional techniques to reduce the query burden. In addition, when features become higher-dimensional, humans may perceive a wider threshold of indistinguishability, which corresponds to a larger threshold of $d^{\text{disc}}$. This phenomenon was also observed in Fig. 5, where increasing the threshold step by step beyond the real human range (Fig. 3) eventually caused the our framework to fail to learn policies. Developing mechanisms that can adaptively handle such broader indistinguishability ranges will be an important direction for future research.
Limitations and Opportunities of VLM-based Preference Annotation: Another important point emerging from our VLM-based experiments is the challenge of relying on VLMs as annotators. Limitations such as weak spatial understanding and overconfidence in discrete choices restrict their current reliability for preference labeling [20, 19]. Our evaluation should therefore be regarded as a basic proof-of-concept, representing an initial step to verify feasibility. While preliminary, these findings highlight both the potential and the current limitations of integrating VLM feedback, and they point to opportunities for future research on more robust methods in PbRL.
Domain Scope and Future Extensions Beyond Quadruped Locomotion: Our experiments have primarily focused on quadruped locomotion tasks, as this domain provides a representative and challenging benchmark for evaluating PbRL frameworks. Quadruped robots require learning diverse gait patterns under high-dimensional dynamics, making them a suitable testbed for studying query efficiency and policy diversity. While the proposed framework is not limited to locomotion and could, in principle, be extended to other robotic domains such as manipulation or humanoid control, a thorough exploration of these directions lies beyond the scope of this paper. We consider this an important avenue for future work, as demonstrating the applicability of our framework across a broader range of skills and platforms would further strengthen its generality.